To increase the oscillation period of a simple pendulum by 41%, how much change is required in its effective length?
1
Expert's answer
2021-09-20T10:00:35-0400

Given:

"T=1.41T_0"


The period of oscillation of a simple pendulum is given by

"T_0=2\pi\sqrt{\frac{l_0}{g}}"

So

"1.41T_0=2\pi\sqrt{\frac{l}{g}}"

"1.41*2\pi\sqrt{\frac{l_0}{g}}=2\pi\sqrt{\frac{l}{g}}"

"1.41\sqrt{l_0}=\sqrt{l}"

Finally,

"l=1.99l_0"

Answer: the length should to be increased by "99\%."